A normally distributed population has a mean of 47 and a standard deviation of 5. What is the probability that a sample of size 7 has a sample mean that is less than 44? complete solution


1
Expert's answer
2021-05-07T10:05:09-0400

"P(\bar X<44)=P(Z<\frac{44-47}{\frac{5}{\sqrt{7}}})=P(Z<-1.59)=0.0559."